the letters in the word BANANAS are to be arranged.A word can be considered formed without being meaningful.Finds event R, Q and T.
R: the words start and end with an A
Q: All the N's in the word are kept together
T: All the A's in the word are kept together

Answer by stanbon(75887)

the letters in the word BANANAS are to be arranged. A word can be considered formed without being meaningful.
Note: There are 7 letters; 3 A's ; 1B ; 2 N's ; 1S
Find event R, Q and T.
R: the words start and end with an A
1st letter: 1 way
7th letter: 1 way
2-6th letters: arranged in 5!/(2! = 120/2 = 60 ways
Total arrangements: 1*1*60 = 60
-------------------------------------
Q: All the N's in the word are kept together
Treat the N's as one letter
# of arrangements: 6!/3! = 120 ways
-------------------------------------
T: All the A's in the word are kept together
Treat the A's as one letter
# of arrangements: 5!/2! = 60 ways
==========================================
Cheers,
Stan H.
-----